Problem 7-43 (LO. 7)

Mario, a single taxpayer with two dependent children, has the following items of income and expense during 2019:

Gross receipts from business $144,000
Business expenses 180,000
Net capital gain 22,000
Interest income 3,000
Itemized deductions (state taxes, residence interest, and contributions) 24,000

a. Determine Mario's taxable income or loss for 2019.

If required, use the minus sign to indicate a negative AGI or a loss.

Adjusted gross income $
Less: itemized deductions
Less: Deduction for qualified business income
Loss $

c. Determine Mario's NOL for 2019.

Mario's NOL is $.
Answer #1

Mario's Business income

Gross receipts = $144000

less Business Expenses = -180000

Loss from Business = -36000

AGI would be ,

Business Loss = -36000 + 22000 + 3000 = -11000

AGI = - 11000,

Adjusted gross income -11000
Less: itemized deductions 24000
Less: Deduction for qualified business income 0
Loss -35000  

NOL = business income less business expenses = 144000 - 180000 = 36000 (or -36000)